OpenStudy (anonymous):

write the equation of the parabola with focus (1, -3) and vertex (5, -3)??

OpenStudy (amistre64):

it opens to the left

OpenStudy (amistre64):

the a value = 4 y^2 =16x modify for off centered

OpenStudy (anonymous):

alright thanks, but how did you know that

OpenStudy (amistre64):

opps; opens left

OpenStudy (amistre64):

the line between the vertex and foci is straight along the y = -3 axis

OpenStudy (amistre64):

the vertex is the bend and the focus is inside the curve

OpenStudy (amistre64):

the distance from the vertex to the focus is the "a" value in the parabola formula from geometery

OpenStudy (amistre64):

y^2 = -4ax

OpenStudy (anonymous):

is that the directrix?

OpenStudy (amistre64):

the directix is on the other side of the vertex; the distance 'a'

OpenStudy (amistre64):

the directix would be at x=9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome :) the complete answer is: (y+3)^2 = -16(x-5)

OpenStudy (anonymous):

lol thanks again :)

OpenStudy (anonymous):

do you know what the latus rectum of a ellipse is?

OpenStudy (amistre64):

sounds painful; .... i prolly do, but just dont know the proper term for it :)

OpenStudy (anonymous):

haha i was out of school and i come back i was like wtf???

OpenStudy (amistre64):

it is defined as the chord running thru the focus

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

amistre64 come back to mine please

OpenStudy (amistre64):

(y+3)^2 = -16(x-5) plug in x = 1 -16(-4) = y^2 + 9 + 6y = -64 y^2 +6x +73 = 0

OpenStudy (amistre64):

6x shoulda been 6y

OpenStudy (amistre64):

the latus rectum is the line thru the foci and those y values

OpenStudy (anonymous):

but thats a parabola right? the ellipse's equation is (x-3)^2/16 + (y+6)^2/4 = 1

OpenStudy (amistre64):

i forgot how to multiply lol

OpenStudy (amistre64):

same concept; its the line thru the focus that runs perp to the center line

OpenStudy (anonymous):

oh ok well, thank you

OpenStudy (amistre64):

we got about a min b4 they shut the site down for maintainence

OpenStudy (anonymous):

dang